Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
bersanetti, tapia, nardecchia, chiummo, tacca - 23:34 Tuesday 20 July 2021 (52615) Print this report
Relock of the ITF for the working point tuning

The goal of the shift was to relock the ITF up to CARM offset zero, in order to start the ICS/TCS tuning of the ITF working point. Unfortunately this could not be done, as we never got past 300 mW of arm powers:

  • at the beginning of the shift we had quite some hard time in locking the DRMI; we soon found that the IMC was out of place of about 800 um, as FmodErr was badly mistuned; we proceeded to move the MC_F0_Z setpoint but we soon found that the IMC cannot be unlocked anymore with the proper relay (or, equivalently, the VPM button); we moved the IMC anyway, just discharging the thermal correction every couple of steps (FIgure 1);
  • after this, we could relock the DRMI quite easily, after the usual pre-alignment; at this point we started to suffer from several unlocks, all of them around 200 mW of arm powers; basically all of them were caused by the 9-10 Hz oscillation (Figures 2 to 6, UTCs in the filename);
  • during these tests, we observed that the DARM gain seemed too low once the DOF is controlled with the B1p_56MHz_I_ARM_NORM signal; we saw, looking at the DARM UGF line and also the DPHI monitor, that the demodulation phase was quite mistuned (~1 rad); we proceeded to re-tune it at every step of the CARM offset reduction, as this has been seen to change; however, in the last (and best) trial we made, it was more or less optimal at 3.85 all the way (this is default in the locking script); to be rechecked in next attempts, especially at 300 mW or arm powers; the same mistuning was not apparently present in the QPD signals from the same bench;
  • later the SDB2_SBE vertical control opened, and it was quickly restored; the cause of this is not clear yet;
  • in the last trial, things seemed to work better keeping a higher gain for the SRCL loop w.r.t. previous attempts (manually, 0.5 @ 100 mW, 0.6 @ 200 mW, 0.8 @ 300 mW); given there is no statistics in this, it's hard to conclude this was effective, but it should be checked.

We leave the arms locked on the IR beam.

Images attached to this report
Comments to this report:
chiummo, Bersanetti, Tacca - 9:04 Wednesday 21 July 2021 (52618) Print this report

For the record, the software button to unlock the IMC does not work since at least aug '20. Yesterday we checked that the software relay works as intended, that the DAC channel (DAC 26 in rack 2 of EEroom - channel 6)  is responding as expected switching from 0V to -5V when the status of IMC_unlock is changed and that the cable is connected from the DAC to IMC_Fr_Cr input of the rampeauto of the IMC.
Fmoderr tuning can be performed safely even without unlocking the IMC, but it is indeed convenient to have the possibility to remotely unlcok the IMC, so we will investigate further to restore this functionality.


tacca - 11:04 Wednesday 21 July 2021 (52619) Print this report
The temperature in the DET lab changed a lot in the last couple of days. The blades of the vertical filters of the mSASes are very sensitive to this kind of temperature variations and the vertical position changes accordingly. Therefore, when there are large temperature variations (as when the tower is vented), the vertical position is checked and readjusted. In the particular case of yesterday evening the vertical actuator reached its high limit and the loop opened.
In the attached plot, the trend of the temperature in the lab and on the F1 of SDB2 mSAS.

Please, try to keep the temperature inside the DET lab as constant as possible.
Images attached to this comment
derossi, spinicelli, chiummo, coulon - 16:01 Friday 23 July 2021 (52646) Print this report

We switched the IMC rampeauto from AUTO to MANUAL mode and the IMC Unlock button is now working. (To be noticed that for the IMC we do not send the ramp but the natural movement of the mirror is sufficient, for this reason the auto mode is not needed).

Search Help